At which vertex is the objective function C = 3x - 4y maximized?
Answer by drk(1908)

First, we solve for y as
y = (3/4)x - C/4
If x = 0, then y = C/4
If y = 0, then x = C/3.
--
C = 3x - 4y.
C(0,C/4) = 3*0 - 4(C/4) = -4
C(C/3,0) = 3*(C/3) - 4*0 = C
--
Now it gets tricky.
If C > -4 then (C/3) is a max.
Otherwise, C/4 is the max.